I sent this one before but made some misstakes sorry. In the question u is for population mean and o is for Population standards deviation thank you .
Let X be a normally distributed random variable with u=100 and o=15 . The probability that is between 80 and 120 is (to the nearest whole percent)
a. 8 % b. 82 % c. 91% d. answer not listed

Answer by stanbon(75887)

In the question u is for population mean and o is for Population standards deviation thank you .
Let X be a normally distributed random variable with u=100 and o=15 The probability that is between 80 and 120 is (to the nearest whole percent)
a. 8 % b. 82 % c. 91% d. answer not listed
--------------
If you have a Ti calulator use normalcdf(80,120,100,15) = 0.8175...
-----
If you do not find the z-scores of 80 and of 120:
z(80)=(80-100)/15 = -1.333333
z(120)=(120-100)/15 = 1.333333
Probability you want is normalcdf(-1.33333,1,33333)=0.8175...
==================
Cheers,
Stan H.